OpenStudy (anonymous):

need help determining what equation to use What is distance around the graph of r = 6 cos θ ?

OpenStudy (anonymous):

The arc length formula doesnt fit because there are no bounds

OpenStudy (irishboy123):

if you faff around with various **convenient** values for \(\theta \) you'd get a feel that you need to run through \( 0 \to 2 \pi\) IOW sketch it out!! or use a computer to draw it..... it's a circle of diameter 6. so you can say that \(C = \pi D\). which i guess is arc length formula simpliciter :p

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

wouldnt the diameter be 12 though? with the length around the circle being 12 pi

OpenStudy (irishboy123):

just plot it and see. it's as boring as \(\theta = 0\) so r = 6 if we go \(\theta = \pi /2\) we have r = 0 so we are at the origin and we might expect that \(\theta = \pi \), we are back where we started cos r = -6 then fill in the in between bits, eg \(\theta = \pi/4\). or switch it back into Cartesian is often a good trick if you can see it better from there. in this case you just need to complete a square! or use desmos https://www.desmos.com/calculator

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

speaking of converting to Cartesian . How would I do that for this http://prntscr.com/ai9hsf